Problem 24.57 Part A Manufacturing Plant Power A manufacturing plant uses 2.51 kW of electric power provided by a 80.0 Hz ac generator with an rms voltage of 400 V . The plant uses this power to run a number of high-inductance electric motors. The planrs total resistance is R 29.0 Ω and its inductive reactance is XL 40.0 Ω What is the total impedance of the plant? Submit My Answers Gve Up Part B What is the plants power factor? Submit My Answere Give Up Part C What is the rms current used by the plant? Submit Part D What capacitance, connected in series with the power line, willincrease the plants power factor to unity? AF Submit My Answere Give Up Part E the power factor is unity, how much current is needed to provide the 251 kW of power needed by the plant? Submit My Anawere Give Up
Answer #1

(A) impedance = sqrt(40^2 + 29^2) = 49.4 ohm


(B) POwer factor = R / Z = 0.725

(C) Irms = Vrms / Z = 460 / 49.4 = 9.31 A


(d) R / Z = 1

R= z

Xc= 40 ohm

40 = 1 / (2 x pi x 60 x C)

C =66.3 x 10^-6 F = 66.3 uF

(e) P = V I

2.51 x 10^3 =460 I

I = 5.46 A